# Extracting Signal Electron Trajectories in the COMET Phase-I Cylindrical Drift Chamber Using Deep Learning

### v1

@@ 6.4 Sensitivity to Scaling Methods
Another critical aspect of our investigation was the sensitivity of the model to different scaling methods, as illustrated in Figure 16. We compared two scaling techniques: the limited scaling and the linear scaling. The limited scaling method was designed to double the resolution for critical ranges of hit time and ADC compared to the linear scaling as discussed in Sec.3.5.2.
@@
To further investigate, we analyzed the performance with low-bit precision inputs. As expected, with lower precision such as 2-bit and 4-bit, the limited scaling method outperformed the linear scaling. As illustrated in Figure 17, the 2-bit case shows that the limited scaling results in finer resolution in the critical ranges of hit time and ADC, highlighting its effectiveness in retaining important features in low-bit scenarios.
@@
Figure 17: Comparison of signal and background hit distributions with the annotation of 2-bit input case: (a) using limited scaling and (b) using linear scaling. Both plots show the distribution of hits over scaled hit time and ADC with a log-scaled y-axis. The shaded areas represent different states for low-bit precision inputs, specifically in the case of 2-bit precision. Only three states (”00”, ”01”, ”10”) are illustrated in the graphs; the ”11” state is reserved for no hit and is not displayed here. The limited scaling method in (a) provides finer resolution in signal-dominant areas, enhancing model performance in low-bit precision scenarios.
@@ 6.5 Sensitivity to $\Delta\phi$ Feature
To assess the impact of including the $\Delta\phi$ feature on model performance, we conducted experiments comparing models trained with and without this feature, which corresponds to 3-channel and 2-channel input images in our setup respectively. The $\Delta\phi$ feature represents the azimuthal angle difference between the CTH trigger and each cell, which could potentially enhance the model’s ability to distinguish signal hits from background hits. Our evaluation focused on two primary input image sizes: 128$\times$128 and 288$\times$288. The Purity-Retention curves for these configurations are presented in Figure 18.
@@
The results show that the models trained with and without the $\Delta\phi$ feature exhibit similar performance in terms of Purity and Retention across both input sizes, with slightly better performance when including the $\Delta\phi$ feature. This suggests that the hit time and ADC features are sufficient to classify the signal and background hits effectively, and the addition of the $\Delta\phi$ feature does not provide significant additional discriminatory power.
@@ 7 Conclusion
    Broader Context and Future Directions: While this paper focuses on signal track extraction in the COMET experiment, it represents only a part of the broader range of the COMET analysis. The ultimate goal of the COMET experiment extends beyond tracking, aiming to determine the energy of electrons from signal hits. Our future work will continue to explore and refine deep learning approaches to meet this objective, including choosing appropriate backbone models and other hyperparameter tuning. Initially, we will focus on developing robust models with simulation data by incorporating more realistic background hits. Following this, we will address the gap between simulation and real data to ensure our methods are robust and reliable in practical applications.

### v2

@@ 6.4 Sensitivity to Scaling Methods
Another critical aspect of our investigation was the sensitivity of the model to different scaling methods, as illustrated in Figure 16. We compared two scaling techniques: the limited scaling and the linear scaling. The limited scaling method was designed to double the resolution for critical ranges of hit timing and ADC compared to the linear scaling as discussed in Sec.3.5.2.
@@
To further investigate, we analyzed the performance with low-bit precision inputs. As expected, with lower precision such as 2-bit and 4-bit, the limited scaling method outperformed the linear scaling. As illustrated in Figure 17, the 2-bit case shows that the limited scaling results in finer resolution in the critical ranges of hit timing and ADC, highlighting its effectiveness in retaining important features in low-bit scenarios.
@@
Figure 17: Comparison of signal and background hit distributions with the annotation of 2-bit input case: (a) using limited scaling and (b) using linear scaling. Both plots show the distribution of hits over scaled hit timing and ADC with a log-scaled y-axis. The shaded areas represent different states for low-bit precision inputs, specifically in the case of 2-bit precision. Only three states (”00”, ”01”, ”10”) are illustrated in the graphs; the ”11” state is reserved for no hit and is not displayed here. The limited scaling method in (a) provides finer resolution in signal-dominant areas, enhancing model performance in low-bit precision scenarios.
@@ 6.5 Sensitivity to $\Delta\phi$ Feature
To assess the impact of including the $\Delta\phi$ feature on model performance, we conducted experiments comparing models trained with and without this feature, which corresponds to 3-channel and 2-channel input images in our setup respectively. The $\Delta\phi$ feature represents the azimuthal angle difference between the CTH trigger counter and each cell, which could potentially enhance the model’s ability to distinguish signal hits from background hits. Our evaluation focused on two primary input image sizes: 128$\times$128 and 288$\times$288. The Purity-Retention curves for these configurations are presented in Figure 18.
@@
The results show that the models trained with and without the $\Delta\phi$ feature exhibit similar performance in terms of Purity and Retention across both input sizes, with slightly better performance when including the $\Delta\phi$ feature. This suggests that the hit timing and ADC features are sufficient to classify the signal and background hits effectively, and the addition of the $\Delta\phi$ feature does not provide significant additional discriminatory power.
@@ 7 Conclusion
    Broader Context and Future Directions: While this paper focuses on signal track extraction in the COMET experiment, it represents only a part of the broader range of the COMET analysis. The ultimate goal of the COMET experiment extends beyond tracking, aiming to determine the momentum of electrons from signal hits. Our future work will continue to explore and refine deep learning approaches to meet this objective, including choosing appropriate backbone models and other hyperparameter tuning. Initially, we will focus on developing robust models with simulation data by incorporating more realistic background hits. Following this, we will address the gap between simulation and real data to ensure our methods are robust and reliable in practical applications.